# Counterfactual Effect Generalization: A Combinatorial Definition

### v4

@@ Abstract
The widely used ’Counterfactual’ definition of Causal Effects was derived for unbiasedness and accuracy - and not generalizability. We propose a Combinatorial definition for the External Validity (EV) of intervention effects. We first define the concept of an effect observation ’background’. We then formulate conditions for effect generalization based on their sets of (observable and unobservable) backgrounds. This reveals two limits for effect generalization: (1) when effects are observed under all their enumerable backgrounds, or, (2) when backgrounds have become sufficiently randomized. We use the resulting combinatorial framework to re-examine several issues in the original counterfactual formulation: out-of-sample validity, concurrent estimation of multiple effects, bias-variance tradeoffs, statistical power, and connections to current predictive and explaining techniques.
Methodologically, the definitions also allow us to also replace the parametric estimation problems that followed the counterfactual definition by combinatorial enumeration and randomization problems in non-experimental samples. We use this non-parametric framework to demonstrate (External Validity, Unconfoundness and Precision) tradeoffs in the performance of popular supervised, explaining, and causal-effect estimators. We demonstrate the approach also allows for the use of these methods in non-i.i.d. samples. The COVID19 pandemic highlighted the need for learning solutions to provide predictions in severally incomplete samples. We demonstrate applications in this pressing problem.
@@ 1 Introduction
Suppose a sample has a Data Generating Process (DGP) which we can only partially observe. We would like to know under what conditions effects (e.g., of an agent’s action or an administered treatment) observed in this sample can be taken as true representations of the effect of its variables. The main challenge is that of confounding of effect estimates. Confounding refers to sampling conditions where uncontrolled variation of out-of-sample or in-sample factors bias in-sample effect estimates. Under these conditions, sample effect observations (i.e., difference in outcomes among sample units) reflect not only the true effect of sample factors, but also of their uncontrolled and unobserved factors. Under a broad class of sampling conditions, effects become, in fact, fully unidentifiable. The ability of estimates to generalize across out-of-sample conditions is often described as a desirable property, but it is not always clear what is meant, and optimized for, in these cases. We study generalization, in particular, from the perspective of Experimental Design, and in scenarios with time-varying treatment exposure and confounding. Formulation of specific sampling conditions is the starting point of many theories of causation (Sect.1.1 Effect estimates and Unobserved Variation), but are often abandoned in lieu of complex regression and parametric methods. This last step makes it difficult to formulate crucial statistical properties for effect estimates such as their precision, biases and out-of-sample generalizability. These requirements have been barely established for popular counterfactual-based effect estimation methods, like Shapley value black-box regressions  [Burkart and Huber, 2021, Lundberg and Lee, 2017, Aas et al., 2021] and g-formula  [Chatton et al., 2020], or most supervised prediction methods. In practice, this makes them difficult to use, and for users to determine when their outputs can be trusted (does the sample have enough observations? variables?). We provide definitions and demonstrate how to select subsamples with simultaneous high generalizability and low confoundness. We demonstrate uses when estimating causal effects non-parametrically, or understanding the performance of out-of-sample supervised predictors and black-box explainers. The latter is currently an intensely studied subject in Artificial Intelligence and Machine Learning  [Lundberg and Lee, 2017, Burkart and Huber, 2021]. This article continues progress on the challenge of estimating causal effects non-parametrically, exclusively from effect observations  [F. Ribeiro et al., 2022].
@@
Partition the set $Z$ of DGP variables into two subsets, $Z=X_{m}\cup U_{q}$, the first with $m$ observed variables $X_{m}=\{-1,+1\}^{m}$, and the second with $q$ unobserved $U_{q}=\{-1,+1\}^{q}$. Each sample unit is, consequently, characterized by a subset of observed factors, $x\in X_{m}$. Imagine also the DGP is constrained by precedence relationships among its variables, described by two partially ordered set (posets) $G(Z)=\{G(X),G(U)\}$. The distinct ways in which $\{G(X),G(U)\}$ can be combined is described by the distinct linear extensions among these posets. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}\;\}$, for example, the extensions are
@@
$\mathbf{ab}u_{1}u_{2}u_{3}\,(*),\,\mathbf{a}u_{1}\mathbf{b}u_{2}u_{3},\allowbreak\,\mathbf{a}u_{1}u_{2}\mathbf{b}u_{3},\allowbreak\,\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b},\allowbreak u_{1}\mathbf{ab}u_{2}u_{3}\,(*),\,u_{1}\mathbf{a}u_{2}\mathbf{b}u_{3},\,u_{1}\mathbf{a}u_{2}u_{3}\mathbf{b},\,u_{1}u_{2}\mathbf{ab}u_{3}\,(*),\,u_{1}u_{2}\mathbf{a}u_{3}\mathbf{b},\,\allowbreak u_{1}u_{2}u_{3}\mathbf{ab}\,(*),\,\allowbreak\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b}.$
The different extensions can be seen as different confounding conditions under which sample $X$ can be observed. We imagine real-world samples can be collected under any such conditions, and we would like to estimate the statistical properties of effects across partially observed samples, $U\neq\emptyset$. The likelihood of observation under a $U{-}X$ extension follow, in turn, relationships between the distributions of factors in $X$ and $U$,
@@
where $p_{t}(b\,|\,a)$ and $p_{t}(u_{i}\,|\,a)$ denote the probability of changes in the value of binary factors $a\in X$ and $u_{i}\subseteq U$, given the observation of a $a$ at a time $t=0$, and $\leq$ denotes poset comparability. The first condition implies that the partial order $a<b<Z-\{a\}$ will apply with high likelihood in the resulting sample. The condition correspond to the first extension in Example.1. That is, that there are no confounders between subsequent $a$ changes, and observed effects in the sample emulate, from frequency-based constraints, a typical counterfactual estimator (Sect.1.1 Effect estimates and Unobserved Variation). This is, however, only true for specific subsamples in the other two scenarios (in equal proportion across all factors in the second and in rare subsamples in the third). For short, we refer to $U{-}X$ systems described by the first conditions as under fast sampling, and the last as under slow sampling. We define alternative sample conditions for the (1) separability (whether pairs of causes and confounders can be distinguished) using only effect differences (or their variance) in samples, and more restrictive conditions of (2) identifiability of causes. Of particular interest in real-world large-scale data is the condition where $U{-}X$ relationships (’treatment-background’ assignments) stay constant, while (one of these) subsets are allowed to change freely. Under some extensions, treatment-background assignment in $X$ can become random-like, and independent of unobserved confounders. These considerations lead to a theoretical model to describe how combinatorial and ordinal relations in $U{-}X$ systems impact the generalizability and confoundness in samples’ effect observations and the performance of estimators that use them.
@@ 1.0.1 Alternative Orders and Confounding: A Preliminary illustration
Let’s consider the observed effects of sample factors that are either causes or confounders in a sample with a true cause $a$, a confounder $b$ (with effects $y$ completely correlated with its root cause $a$), and a spurious variable $c$, $X=\{a,b,c\}$. Let’s consider, in particular, differences in observed effects of $a$ and $b$ across different values of other factors. We will use Latin square diagrams as aid to illustrate several concepts in this article. Fig.1(a) shows an example with $4$ factors. Each square is associated with a sample unit $x_{0}$, its reference unit, which stratifies a population (other units) across its cells. A square with reference $x_{0}$ correspond to the set of effect observations with singleton factor differences,
@@
where $\mathcal{P}(X)$ is the powerset (set of sets) of $X$ and ${\Delta}y(x-x_{0})$ is an effect observation in outcome-of-interest, $y\in\mathbb{R}$, after an observed intervention on factors $x-x_{0}$. A single square contains therefore $\binom{m}{2}$ effect observations, and a sample (or all its squares) is limited to $2^{m}$ observations. We will say that the square’s diagonals enumerate, for example, observations of effect for a fixed factor across different (observed) backgrounds. The diagonal effect observations for the cause $a$ and confounder $b$ are
$\hfil\textrm{diagonal}\;a\hfil&\hfil\textrm{diagonal}\;b\hfil\\[8.0pt] \Delta{y}(a)&&\hfil\hskip 30.&\Delta{y}(b)&&\\ &\Delta{y}(a\,|\,b){-}\Delta{y}(a){=}0&\hfil\hskip 30.&&\Delta{y}(b\,|\,a){-}\Delta{y}(b){=}\Delta{y}(a)&\\ \Delta{y}(a\,|\,b)&&{=}0,\hfil\hskip 30.&\Delta{y}(b\,|\,a)&&{=}{-}\Delta{y}(a),\\ &\Delta{y}(a\,|\,bc){-}\Delta{y}(a|b){=}0&\hfil\hskip 30.&&\Delta{y}(b\,|\,ac){-}\Delta{y}(b\,|\,a){=}0&\\ \Delta{y}(a\,|\,bc)&&\hfil\hskip 30.&\Delta{y}(b\,|\,ac)&&\\ \end{array}$
We see that when the cause $a$ is in the sample, it sustains its effects through arbitrary sample imputations (indicated by the conditional statements), while the confounder also assumes negative values. In a square and fully varying set $X$ of factors, confounders change their effect observations several times, while causes remain invariant throughout all enumerated variation.
More importantly, however, this indicates that, to identify confounders in fully-observed samples ($U=\emptyset$), we need samples where we, minimally, observe the effect of both adding factor $a$ to units with $b$, and of adding factor $b$ to units with $a$ (i.e., under alternative orders). These conditions are necessary because $b$ has no effects only in the latter case, $\Delta y(a)+\Delta y(a\,|\,b)>\Delta y(b)+\Delta y(b\,|\,a)$. Samples that include these conditions allows us to see, from effect observations alone, that $b$’s effects were brought by $b$’s correlation with $a$, and not its own independent effect on $y$. This constitutes a simple sample requirement for cases with $m=2$ and $q=0$. The rest of this article generalizes this to the $m$ factor case and the possibility of $q$ unobserved factors.
@@
[…]
According to the principle, under some ordinal conditions on $U{-}X$ factors, effect observations for confounders have an expected effect variance of $0.25\times\rho_{ab}$ (illustrated by the number of times the confounder appears before its root cause in square rows) and causal variables 00 of the true effect $\Delta y(a)$. One such condition is of a fully observed sample, $U=\emptyset$, when the principle allow the non-parametric identification of causes. In this case, the principle is associated with the observation of each effect under the cyclic permutation of all others factors (as illustrated by square rows), and it says that effect variances, ${\rm I\kern-1.99997ptE}[\,\textrm{Var}[\Delta\hat{y}(a)]\,]$, are a sufficient statistic, for a fixed level of cause-confounder correlation. Other conditions, established below, are required otherwise. These conditions correspond, non-parametrically, to the observation of multiple squares.
1.1 Effect estimates and Unobserved Variation
Before formulating necessary conditions for the generalizability and confoudness of effect observation in samples, we give our working definitions of effects and their generalizability. Subsequent sections justify, with increasing mathematical rigor, these combinatorial definitions. According to the Potential Outcomes framework’ definitions for what is a causal effect  [Morgan and Winship, 2007, Rubin, 2005], if $y$ is an outcome of interest and $a$ is a treatment indicator, then the causal effect of $a$ is the difference
@@ 1.1 Effect estimates and Unobserved Variation
where $y_{i}^{\texttt{+}a}$ is the outcome of individual $i$ under the treatment, and, $y_{i}^{\texttt{-}a}$ without the treatment. The central concept behind Eq. (1) was inspired by experimental estimation: by fixing every factor, other than the treatment, we can declare that the observed difference in outcome was certainly caused by the treatment, and the treatment alone. The definition is an ideal, as it is impossible to observe outcomes for an individual, concurrently, in two different and totally fixed conditions. The theory goes that we may, instead, ’fix’ factors in expectation, and across individuals. If the treated and non-treated subpopulations have the same expected values across all factors then any difference between the groups is due to the treatment, given large enough samples. This can be paraphrased with an independence statement: treatment assignment must be independent on all outcome-relevant factors. This rationale led to the notion of Sample Balance in non-experimental estimation  [Rubin, 2005, Morgan and Winship, 2007] and the objective most current counterfactual causal effect estimators maximize.
@@
where $y_{\leq}$ is the observed outcome $y$ of a sample units with, exclusively, the set of factors before $a$ in the permutation order $\pi$, and $y_{>}$ with factors after $a$. Furthermore, $\Pi_{n}(Z/\{a\})$ is a subset of the full set of permutations of factors $Z$ observable between sequential changes in $a$ values, and is determined by the cycle-structure of permutations over $Z/\{a\}$. Sample ordinal patterns, correlations and periodicity determine $\Pi_{n}$. Difference between units in Eq.(3) constitutes observation of effects of the same factor in distinct backgrounds $\pi\in\Pi_{n}(Z/\{a\})$. Each background fixes a set of factors for their ’treatment’ (before $a$ in $\pi$ or, not containing $a$) and ’control’ (after $a$) for sample units. For a fixed order $\pi$, there are $m$ orders that preserve the relative order of in-sample factors. The rows of a Latin-Square diagram, Fig.1(a), describe a single such order for factors $X$. To see this, choose any factor in a square diagram and follow its subsequent factors cyclically (e.g., $b<c<d$ for $a$). The set of enumerable permutations will be generalized to samples, $\Pi_{n}(U=Z/X)$, in Sect.3 Sampling Treatment-Background Assignments.
At the limit, when all factors in $U=Z/X$ are independent, $\Pi_{n}$ correspond to the full set of permutations of size, $|\Pi_{n}|=q!$. At the same time, when all factors are serially correlated, $\Pi_{n}$ correspond to to only one, $|\Pi_{n}|=1$. Sample correlations and fixations in the cycle structure of permutations can drastically reduce $|\Pi_{n}|$, many times asymptotically and even in very large samples. In the next section, we will use sample ordinal relations to describe this relationship in further detail. We will consider how incompleteness of $\Pi_{n}$ in respect to both $X$ and $U$ affects (often in unknown ways) the generalizability of effect estimates. Based on these conditions, we say an effect is Externally Valid (EV) if it holds under the system $U{-}X$ set of unobserved backgrounds,
@@
EV is here the inverse of the variance (i.e., the precision) of observed effects under all asymptotically observable backgrounds, $\Pi_{n}(Z/\{a\})$. The most important element of these definitions are $\Pi_{n}(Z/\{a\})$, subsets of unobservable permutations in a $\{X,U\}$ when $a$ is fixed. This defines causal effects in a way that is almost opposite to Eq. (1). It calls for effects to be observed under large variation, as opposed to no variation. According to Eq.(3,4), the estimate is, in fact, the one with minimal EV. Underlying these definitions and Principle.1 is the observation that, unlike confounders, causes can be defined fundamentally as ’perfect controls’ [Pearl, 2000]. Similar definitions have been explored in work based on robustness (of models) [Peters et al., 2016, Buehlmann, 2020] (Sect.4 Related Work). The non-parametric definitions here frame the problem, in particular, not as simply the invariance (of effects), but invariance of effects under maximum variance of their ’backgrounds’. Shifting the focus to the later, $\Pi_{n}(U{=}X/Z)$, is a key but challenging part of the proposed solution.
@@
Figure 1: (a) $4{\times}4$ Latin-Square (’square’) as sets of effect observation for sample unit $x_{0}\subseteq X$, (b) $3$ samples with increasing random factors $X$ and their ordinal statistics (interval ticks), $m=\{1,2,4\}$, and alternative sampling of confounding factors $U$ (circles)(top), resulting distribution and percentage of enumerable permutations of effect observations across ordinal variables (bottom and right).
2 Confounding Ordinal Statistics and their Unobserved Gaps

### v5

# Sample Observed Effects: Enumeration, Randomization and Generalization

@@ Abstract
The widely used ’Counterfactual’ definition of Causal Effects was derived for unbiasedness and accuracy - and not generalizability. We propose a Combinatorial definition for the External Validity (EV) of intervention effects. We first define the concept of an effect observation ’background’. We then formulate conditions for effect generalization based on their sets of (observed and unobserved) backgrounds. This reveals two limits for effect generalization: (1) when effects are observed under all their enumerable backgrounds, or, (2) when backgrounds have become sufficiently randomized. We use the resulting combinatorial framework to re-examine several issues in the original counterfactual formulation: out-of-sample validity, concurrent estimation of multiple effects, bias-variance tradeoffs, statistical power, and connections to current predictive and explaining techniques.
Methodologically, the definitions also allow us to replace the parametric estimation problems that followed the counterfactual definition by combinatorial enumeration and randomization problems in non-experimental samples. We use this non-parametric framework to demonstrate (External Validity, Unconfoundness and Precision) tradeoffs in the performance of popular supervised, explaining, and causal-effect estimators. We also illustrate how the approach allows for the use of supervised and explaining methods in non-i.i.d. samples. The COVID19 pandemic highlighted the need for learning solutions to provide predictions in severally incomplete samples. We demonstrate applications in this pressing problem.
@@ 1 Introduction
Non-parametric effect estimators (such as permutation tests  [Hoeffding, 1948] and feature importance attribution  [Burkart and Huber, 2021]) can often only be used under strict assumptions of no unobserved factors. Consider, for example, a sample containing units that have received a treatment (’treatment’ units) and others that did not (’control’). To estimate effects, a simple strategy is to permute treatment assignment among all treatment-control pairs that are similar enough across relevant sample variables  [Morgan and Winship, 2007]. This strategy relies, however, on the assumption that paired units are exchangeable besides their treatment status. The assumption is easily violated if units differ in any unobserved way (even in high-dimensional samples it is hard for researchers to make the claim that there are no overlooked, relevant distinctions). The main challenge is that of unobserved confounding. Confounding refers to sampling conditions where uncontrolled variation of out-of-sample factors bias in-sample estimates. Under these conditions, sample observation of effects, for example, reflect not only the true effect of sample factors, but also of uncontrolled and unobserved factors. In Experimental Design, randomization plays a crucial role in alleviating the effects of unobserved confounding  [Zhang and Zhao, 2023, Kempthorne and Doerfler, 1969]. Randomized Experiments provide a solid foundation, and are still the golden rule, for making causal inferences about the effects of treatments in the Sciences. Non-experimental samples can be seen as experimental samples where randomization did not occur perfectly. Despite these known shortcomings, permutation and Shapley-value based tests  [Burkart and Huber, 2021, Ribeiro, 2022a], continue to be used in practice in non-experimental and partially observed data, with little clarity over the resulting effect estimates’ generalizability and biases. We provide a combinatorial definition for effect generalizability and demonstrate how to select subsamples in non-experimental data that lead to effect estimates with high generalizability and low confoundness. We demonstrate uses of these concepts when estimating causal effects non-parametrically or when explaining the performance of supervised predictors. The latter is currently an intensely studied subject in Machine Learning and Artificial Intelligence  [Lundberg and Lee, 2017, Burkart and Huber, 2021].
@@
A sample’s Data Generating Process (DGP) refers to the underlying mechanism or model that generated the sample. Partition the set $Z$ of DGP variables into two subsets, $Z=X_{m}\cup U_{q}$, the first subset with its $m$ observed variables $X_{m}=\{-1,+1\}^{m}$, and the second with $q$ unobserved $U_{q}=\{-1,+1\}^{q}$. Imagine also variables in $X$ and $U$ are constrained by precedence relationships, as described by two partially ordered set (posets) $G(Z)=\{G(X),G(U)\}$. The distinct ways in which $\{G(X),G(U)\}$ can be combined is described by the distinct linear extensions among these posets. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}\;\}$, for example, the extensions are
@@
$\mathbf{ab}u_{1}u_{2}u_{3}\,\,\mathbf{a}u_{1}\mathbf{b}u_{2}u_{3},\allowbreak\,\mathbf{a}u_{1}u_{2}\mathbf{b}u_{3},\allowbreak\,\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b},\allowbreak u_{1}\mathbf{ab}u_{2}u_{3}\,\,u_{1}\mathbf{a}u_{2}\mathbf{b}u_{3},\,u_{1}\mathbf{a}u_{2}u_{3}\mathbf{b},\,u_{1}u_{2}\mathbf{ab}u_{3}\,\,u_{1}u_{2}\mathbf{a}u_{3}\mathbf{b},\,\allowbreak u_{1}u_{2}u_{3}\mathbf{ab},\,\allowbreak\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b}.$
The different extensions can be seen as different confounding conditions, and orders, under which sample $X$ can be observed. We imagine real-world samples can be collected under any such conditions, and we would like to formulate statistical properties for effects across partially observed samples, $q>0$. The likelihood of a $U{-}X$ extension follow, in turn, relationships between the DGP probability distributions for $X$ and $U$,
@@
[…]
@@ 1.0.1 Alternative Orders and Confounding: A Preliminary illustration
For illustration, consider the observed effects of sample factors in a sample with a true cause $a$, a confounder $b$ (with effects on $y$ fully correlated with its root cause $a$), and a spurious variable $c$, $X=\{a,b,c\}$. We will use Latin square diagrams (’squares’) as aid to illustrate several concepts in this article. Fig.1(a) shows a $4$-factor square. Each square is associated with a sample unit $x_{0}$, its reference unit, which stratifies the sample (i.e., other units) across its cells. A square with reference $x_{0}$ correspond to the set of effect observations with singleton factor differences,
@@
where $\mathcal{P}(X)$ is the powerset (set of sets) of $X$ and ${\Delta}y(x-x_{0})$ is an observation of effect in an outcome-of-interest, $y\in\mathbb{R}$, after an intervention on factors $x-x_{0}$. A complete square contains therefore $\binom{m}{2}$ effect observations, and a sample (or all its squares) is limited to $2^{m}$ observations. We will say that the square’s diagonals enumerate, for example, observations of effect for fixed factors across different (observed) ’backgrounds’. For the cause $a$ and confounder $b$, effect observations are therefore
[…]
We see that when the cause $a$ is in the sample, it sustains its effects through arbitrary sample imputations (indicated by the conditional statements), while the confounder also assumes negative values. In a square with $X$ factors, observed under different orders, effect observations for confounders change many times, while causes remain invariant throughout the observed range of background variation.
More importantly, however, this indicates that, to identify confounders we need samples where we, minimally, observe the effect of both adding factor $a$ to units with $b$, and of adding factor $b$ to units with $a$ (i.e., under alternative orders). These conditions are necessary because $b$ has no effects only in the latter case, $\Delta y(a)+\Delta y(a\,|\,b)>\Delta y(b)+\Delta y(b\,|\,a)$. Samples that include these conditions allows us to see, from effect observations alone, that $b$’s effects were brought by $b$’s correlation with $a$, and not its own independent effect on $y$.
@@
For any factors $a,b\in Z$, observing their effects under conditions $\Delta y(a\,|\,b)$ and $\Delta y(b\,|\,a)$, or in ordinal relation $a>b$ and $b>a$, is necessary to identify, from effect observations, whether $a$ confounds $b$ (and vice versa). When observed under such conditions, effect variances, $\textrm{Var}[\Delta\hat{y}(b)]$, for causes, confounders and spurious factors become distinct, and have expected values 00, $\nicefrac{{\rho_{ab}}}{{4}}\Delta y(a)$, $\nicefrac{{1}}{{2}}\Delta y(a)$, where $\rho_{ab}$ is the Spearman’s rank correlation coefficient between confounder and its root cause, and $\Delta y(a)$ the root cause’s effect.
According to the principle, under specific ordinal conditions on $U{-}X$ factors, effect observations for confounders have an expected effect variance of $0.25\times\rho_{ab}$ (i.e., the number of times the confounder appears before its root cause in square rows) and causal variables 00 of the true effect $\Delta y(a)$. One such condition is of a completely observed sample, $q=0$. In this case, the principle is associated with the observation of each effect under the cyclic permutation of all others factors (as illustrated by distinct square rows). The principle then states that effect variances, ${\rm I\kern-1.99997ptE}[\,\textrm{Var}[\Delta\hat{y}(a)]\,]$, are then a sufficient statistic to identify confounders under a fixed level of cause-confounder correlation $\rho_{ab}$. Other conditions, $q>0$, are established below.
1.1 Effect estimates and Unobserved Sample Variation
Before formulating necessary conditions for the generalizability and confoudness of effect observation in generic samples, we formulate our working definition of effects, their generalizability, and a roadmap to the article. According to the Potential Outcomes framework’ definitions for what is a causal effect  [Morgan and Winship, 2007, Rubin, 2005], if $y$ is an outcome of interest and $a$ is a treatment indicator, then the causal effect of $a$ is the difference
@@ 1.1 Effect estimates and Unobserved Sample Variation
where $y_{i}^{\texttt{+}a}$ is the outcome of individual $i$ under the treatment, and, $y_{i}^{\texttt{-}a}$ without the treatment, $a\in X$. The central concept behind Eq. (1) was inspired by experimental estimation: by fixing every factor, other than the treatment, we can declare that the observed difference in outcome was certainly caused by the treatment, and the treatment alone. The definition is an ideal, as it is impossible to observe outcomes for an individual, concurrently, in two different and totally fixed conditions. The theory goes that we may, instead, ’fix’ factors in expectation, and across individuals. If the treated and non-treated subpopulations have the same expected values across all factors then any difference between the groups is due to the treatment, given large enough samples. This can be paraphrased with an independence statement: treatment assignment must be independent on all outcome-relevant factors. This rationale led to the notion of a sample’s ’balance’ in non-experimental estimation  [Rubin, 2005, Morgan and Winship, 2007] and the objective most current counterfactual causal effect estimators maximize.
@@
where $y_{\leq}$ is the observed outcome $y$ of a sample unit with, exclusively, the set of factors before $a$ in the permutation order $\pi$, and $y_{>}$ of units also containing the factors after $a$. Difference between units in Eq.(3) constitutes observations of effects of the same factor in distinct backgrounds, $\pi\in\Pi_{n}(Z/\{a\})$. Each background fixes a set of factors for their ’treatment’ (before $a$ in $\pi$ or, similarly, sample effects that do not reflect the effect of $a$) and ’control’ (after $a$) for sample units.
Furthermore, $\Pi_{n}(Z/\{a\})$ denotes a subset of the $(m+q)!$ full set of permutations of factors $Z$ observable between sequential changes in $a$ values. What permutation subsets can offer guarantees over the generalizability of sample effect estimates? We formulate the concept of External Validity increasing (EV-increasing) permutations as those that can increase the generalizability of effects in samples. The article’s main argument is structured as follows,
Section Summary
2. Ordinal Statistics and Sample Confounding Gaps formulates the concept of sample ’gaps’ to describe the action of confounders $U$ across in-sample factors $X$.
3. Sampling Treatment-Background Assignments formulates permutation subsets $\Pi_{n}(Z/\{a\})$ that are EV-increasing and CF non-increasing, and their relationship to sample gap statistics.
4. Square Enumeration and Gap Randomization Tests formulates non-parametric tests for effect background randomization based on the previous concepts.
5. Related Work recasts the approach under graphical and structural causal frameworks and relate it to other non-parametric effect estimators.
6. Sample Sizes formulates resulting sample size requirements for effect generalizability.
Based on these concepts, we say an effect is Externally Valid (EV) if it holds across the subset $\Pi_{n}(Z/\{a\})$ of backgrounds,
@@
EV is here the inverse of the variance (i.e., the precision) of observed effects under all (e.g., asymptotically) EV-increasing backgrounds, $\Pi_{n}(Z/\{a\})$, in an $U-X$ system for a treatment $a$. This defines effects in a way that is almost opposite to Eq. (1). It calls for effects to be observed under large variation, as opposed to no variation. According to Eq.(3,4), the counterfactual ideal in Eq.(1) is, in fact, the estimate with minimal EV. Underlying these definitions and Principle (1) is the observation that, unlike confounders, causes can be fundamentally defined as ’perfect controls’  [Pearl, 2000]. Similar definitions have been explored in model robustness research  [Peters et al., 2016, Buehlmann, 2020]. The non-parametric definitions here frame the problem, in particular, not as simply the invariance (of effects), but invariance of effects under maximum variance of their ’backgrounds’. Shifting the focus to the later, $\Pi_{n}(Z/\{a\})$, is a key but challenging part of the proposed solution.
@@
Figure 1: (a) $4{\times}4$ Latin-Square (’square’) as sets of effect observation for a sample unit $x_{0}\subseteq X$, (b) $3$ samples (horizontal lines, top) with increasing dimension $m=\{1,2,4\}$ and their ordinal statistics (interval ticks) , as well as alternative samples of confounding unobserved factors $U$ (circles)(top), resulting distribution and percentage of enumerable permutations of effect observations across ordinal variables (bottom and right).
2 Ordinal Statistics and Sample Confounding Gaps